OpenStudy (anonymous):

i need help with solving polynomialequations like so: (d-7)(d+4/3)=0

OpenStudy (anonymous):

(a+b)(c+d)=ac+ad+bc+bd

OpenStudy (anonymous):

but d is the same so how would d=c?

OpenStudy (anonymous):

if d is a how could it be C?

OpenStudy (anonymous):

d={7, -4/3}

OpenStudy (anonymous):

do i distribute?

OpenStudy (anonymous):

no, it's already factored

OpenStudy (anonymous):

hmmm, ok, like this: (x+y)(u+v)=xu+xv+yu+yv apply this into your equation x=d, y=-7, u=d, v=4/3

OpenStudy (anonymous):

no way ttt, d can't equal -7, that makes the equality false!

OpenStudy (anonymous):

im solving it it will =0 like this, (x-4)(X+2)=0 so x-4=0 and x+2=0

OpenStudy (anonymous):

so x=4 or x=2

OpenStudy (anonymous):

but i dont get it

OpenStudy (anonymous):

what don't you get?

OpenStudy (anonymous):

oh dear =)) d=7 d=-4/3 something wrong with my head just now lol

OpenStudy (anonymous):

howdoes x=4 when it is x-4=0

OpenStudy (anonymous):

if you add 4 to both sides of the equation x - 4 = 0 you get x = 4

OpenStudy (anonymous):

so for my first equation d-7=0 and d+4/3=0?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

what do i do from there i add -7 and 4/3 to each side?

OpenStudy (anonymous):

no, you add 7 and subtract 4/3

OpenStudy (anonymous):

thats what i meant thanks alot for your help tho

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

what would my answer be if they both =0 would it be d=0?

OpenStudy (anonymous):

they both don't equal 0, only one at a time

OpenStudy (anonymous):

wait so i do 4/3-d or 0

OpenStudy (anonymous):

or 4/3-0*

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what would i do -7+d or -7+0???

OpenStudy (anonymous):

what? d is two distinct answers d ={7, -4/3}

OpenStudy (anonymous):

but if d-7=0 and d+4/3=0 what would i do then

OpenStudy (anonymous):

hi people :P

OpenStudy (anonymous):

samarrel?

OpenStudy (anonymous):

that means that d has two solutions 7 and 4/3

OpenStudy (anonymous):

hey people who wants to play uno

OpenStudy (anonymous):

so thats my answer?

OpenStudy (anonymous):

sam?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

(d-7)(d+4/3)=0 d-7=0, d=7 because 7-7=0 d+4/3=0, d=-4/3 because -4/3+4/3=0 which part don't you understand? dont make it too complicated.

OpenStudy (anonymous):

wouldnt it be 7+7?

OpenStudy (anonymous):

why would it be 7+7? d=7, so you change letter d in the equation into number 7. you didn't touch the minus sign at all so how can it become +

OpenStudy (anonymous):

because d-7=0 add 7 to each side right?

OpenStudy (anonymous):

ohhh ok i get it if it = 0 than it would be corect right?

OpenStudy (anonymous):

that's what people always do, I dont. why add more stuff? think this way. you have 7 on the right, now change it to the left. everytime you change side of anything, you change the sign. that number 7 on the left, you have - before it, so it's -7. now you move -7 to the left, you change the sign. So -7 on the right become +7 on the left. done

OpenStudy (anonymous):

now try this, a-6+12=2, lets see if you get it right.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

idk do i add 6 to each side?

OpenStudy (anonymous):

no no, dont add anything, move it to the other side. what you have on the left? you have (-6) and (+12) right? so change the side. remember, CHANGE SIDE, CHANGE SIGN. (-6) from left to right become +6, +12 left to right be come -12. can u do the rest now?

OpenStudy (anonymous):

a-12=2+6?

OpenStudy (anonymous):

that 12 there, why is it still on the left. It should be changed to the right to become -12. a=2+6-12 like that try another one x+5-7=1 remember, if you change its side, also change its sign, if not, nothing happen

OpenStudy (anonymous):

x=1-5+7?

OpenStudy (anonymous):

absolutely right. how about 5-9=7-x

OpenStudy (anonymous):

-7+5-9=x

OpenStudy (anonymous):

should be -x, because -x wasn't changed side, right? -7+5-9=-x u understand this part?

OpenStudy (anonymous):

ohh ok yea

OpenStudy (anonymous):

so I think you get it now. back to the problem, you have d-7=0 or d+4/3=0 just do the same, change side, change sign.

OpenStudy (anonymous):

0=7+d?

OpenStudy (anonymous):

??? no, keep d in one side, push everything else to the other side, because you're finding d here. do again?

OpenStudy (anonymous):

ohh ok because im ffinding d it would be 0+7=d?

OpenStudy (anonymous):

that's alright, but you just begin with this method, so try follow the rule d-7=0 -7=-d (see, change side, change sign) d=7 (again, change side, change sign)

OpenStudy (anonymous):

if you move anything from left to right or right to left, change the sign that go with it. eg. -7 on the left becomes 7 on the right or x on the right, become -x on the left

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

bla bla bla

OpenStudy (anonymous):

d+4/3=0 ? d=?

OpenStudy (anonymous):

-4/3+0=d

OpenStudy (anonymous):

my aint used to make me do this stuff for the whole day. she's a teacher

OpenStudy (anonymous):

that's ok but you shouldn't do it that way because it could easily get wrong. you have d+4/3=0 what you wanna find should be on the left, so d stay. 4/3 on the left, moved to the right, becomes -4/3 so d=-4/3

OpenStudy (anonymous):

i want to do this stuff all day just dont have the stuff to do it so my friend aarman who is great on this site showed me it so i come on here

OpenStudy (anonymous):

what happend to 0?

OpenStudy (anonymous):

lol, 0 is nothing, just thought it away. anything add up with 0 become itself. right, can you solve this: d-1/4+5=0 remember, d stay, everything else go.

OpenStudy (anonymous):

throw* lol

OpenStudy (anonymous):

d=-1/4-5?

OpenStudy (anonymous):

you forgot to change sign of -1/4. can you see that mistake?

OpenStudy (anonymous):

on the left, it's -1/4, so on the right, what should it be?

OpenStudy (anonymous):

1/4-5

OpenStudy (anonymous):

that's right. I'm gonna go to sleep now, but try this last one first. -4d+5-3+12=1-7d+2/3 remember bring all the d to the left and all the number to the right. and do slowly step by step, no rush

OpenStudy (anonymous):

1+4-5+3-12=d

OpenStudy (anonymous):

not really, you know left and right right? (left)=(right) look at the term d first, we have -4d on the left, so keep it, -7d on the right, bring it to the left, change the sign, become 7d. so on the left you have -4d+7d on the right, there's 1 and 2/3, keep them. there're +5-3+12 on the left, bring them to the right, change sign, become -5+3-12 so the right: 1+2/3-5+3-12 then equation is -4d+7d=1+2/3-5+3-12 compare these two: -4d+5-3+12=1-7d+2/3 (origin) -4d+7d=1+2/3-5+3-12 (after changing) I'll put some brackets for clarity -(4d) + 5 -3 +12 = 1 -(7d) +2/3 (origin) -(4d) +(7d) = 1 +2/3 -5 +3 -12 (after changing)

OpenStudy (anonymous):

ok man, I'm gonna go, but this will help you http://www.mathexpression.com/solving-linear-equation.html watch the video :)